# Covariant Measures of Non-Markovianity in Curved Spacetime

### v1

@@ III.1 Operational Definition and Resource-Theoretic Perspectives
where the supremum ranges over causal testers (normalized multi-time instruments obeying the usual recursion $\operatorname{Tr}_{\mathrm{out}_{k}}T^{(k)}=T^{(k-1)}$ with $T^{(0)}=\mathbb{I}/d$), and the ancilla dimension can be taken no larger than the product of input dimensions up to slot n1n{-}1. Operationally, (21) is the maximal distinguishing advantage of an optimal entanglement-assisted, adaptive probing strategy between $\Upsilon$ and its closest Markovian approximation; faithfulness follows immediately since $N(\Upsilon)=0$ iff $\Upsilon\in\mathcal{M}$ [22, 43].
@@
The distance $N$ defined in (20)–(21) has the basic features expected of an operational resource measure for memory. (i) Faithfulness and convexity: $N(\Upsilon)=0$ iff $\Upsilon\in\mathcal{M}$, and $N$ is convex on the affine space of combs: $N(p\Upsilon{+}(1{-}p)\Gamma)\leq p\,N(\Upsilon)+(1{-}p)\,N(\Gamma)$. (ii) Stability (Lipschitz continuity): by the triangle inequality for the strategy norm, $\bigl|N(\Upsilon)-N(\Gamma)\bigr|\leq\|\Upsilon-\Gamma\|_{\rm comb}$, which we use later to relate $N$ to entropic diagnostics. (iii) Monotonicity under Markovian supermaps: if $\mathbb{S}$ is a free transformation obtained by slot-wise pre/post-processing with CPTP channels and memoryless side channels that preserve the comb causality constraints (i.e., $\mathbb{S}[\mathcal{M}]\subseteq\mathcal{M}$), then
@@ III.2 Witnessing and Bounding Non-Divisibility
where $K_{k:j}=\int_{\gamma_{k:j}}\langle H_{\mathrm{int}}(\lambda)\rangle_{E}d\lambda$ (often zero) and the double commutator encodes dissipation via the Wightman function $W$. The Choi operator approximates
@@ III.3 Multi-Time Phenomena and Superactivation in composite systems
Superactivation occurs whenever $\Delta_{\mathrm{act}}>0$, indicating memory effects undetectable in marginal probes but revealed by coherent multi-time interventions (cf. cross-time links in Fig. 2(b)).
@@
where cross-segment kernels $\{\eta_{m-l}\}_{l\in S_{1},m\in S_{2}}$ add coherently. Curvature-induced long-range tails in $W$ (e.g., near horizons) can make intra-segment sums positive (CP maps) while cross terms drive overall negativity, yielding $\Delta_{\mathrm{act}}>0$ and a finite LB (39) [12, 11, 14, 16, 15].
@@ IV Memory kernel structures in curved spacetimes: numerical benchmarks and physical insights
To isolate curvature-induced memory, we employ Hadamard subtraction [27, 30, 51]: the universal singular part of $W$ (arising from the Minkowski-like local structure near coincidence) is removed, leaving the regular, state-dependent remainder that encodes global correlations. This subtraction ensures that $\eta$ reflects genuine non-local effects rather than ultraviolet divergences. Numerically, we discretize the worldline into $n=14$ slots over a total affine parameter range $L=12.0$, with Gaussian switches of width $\sigma=0.4\Delta$ (where $\Delta=L/n$) and use Simpson quadrature with 500 points per integral for accuracy. Geodesic integration for infalling trajectories is handled via solve_ivp with adaptive Runge-Kutta (RK45) stepping, achieving relative tolerances of $10^{-8}$.
@@
Figure 4 displays line cuts of the memory kernel against slot separation together with parameter scans of the kernel’s most negative entry, which we use as a proxy for long–range anticorrelations. Panel 4(a) plots the median magnitude $(\langle|\eta_{i-j}|\rangle)$ versus (|ij|)(|i-j|) on semi–log axes for four worldlines: Minkowski inertial (blue), Minkowski Rindler (orange), Schwarzschild static (green), and Schwarzschild infall (red). Across all scenarios except Schwarzschild–static we observe a dip at |ij|=1|i-j|=1, which arises because adjacent slots correspond to partially overlapping causal diamonds: the convolution of the Wightman kernel with two overlapping window functions produces strong destructive cancellation. As the separation grows the diamonds cease to overlap, this cancellation rapidly weakens, and the median $\langle|\eta_{i-j}|\rangle$ rebounds toward the asymptotic floor set by the intrinsic tail of the correlator (algebraic for inertial/infall, thermal for Rindler/static). The Schwarzschild–static case lacks this feature because proper-time stationarity with a single Tolman thermal scale places the kernel directly on its exponential slope, so the overlap-induced minimum is washed out.
@@
Panels 4(b)–(d) quantify how geometry and kinematics tune the most negative off–diagonal entry. For infall, $(\min\eta)$ becomes more negative as the initial radius $(r_{0})$ increases [Fig. 4(b)], indicating that longer exposure to gentle curvature gradients amplifies the negative lobe at intermediate lags. In the Rindler case, $(\min\eta)$ plateaus for small accelerations and then decreases approximately linearly beyond a knee at $(a\Delta\tau=\mathcal{O}(1))$ [Fig. 4(c)]: below the knee the Unruh correlation time $(\xi_{U}\propto 1/a)$ exceeds the slot width and the kernel is effectively inertial; once resolved, thermal anticorrelations deepen with (a). For static Schwarzschild, $(\min\eta)$ shows a shallow shoulder at small radii before a gentle outward decrease [Fig. 4(d)]. This shape reflects the competition between Tolman heating, which shortens $(\xi(r))$ and strengthens anticorrelations near the horizon, and redshift of the proper-time step $(\Delta\tau=\sqrt{1-2M/r},\Delta t)$ at fixed coordinate sampling, which reduces adjacent–slot overlap and partially offsets thermal screening. Far from the horizon the local temperature drops, the correlation length grows, and $(\min\eta)$ relaxes toward the cooler, nearly flat–space limit.
@@
Figure 5 gathers these kernel-level observations into fully operational statements about the process tensor. Panel 5(a) shows, for $n\leq 5$, the SDP-derived lower-bound proxy for $N(\Upsilon)$ obtained by fitting the tester data to the nearest CP-divisible product of single-step channels (cf. Appendix D). The bound is quantified by the summed absolute prediction error $\sum_{\text{testers}}|\Delta|$, which lower-bounds the strategy-norm distance between the true process tensor and its best Markovian approximation within this restricted class. For our benchmark parameters, the Minkowski inertial worldline yields the smallest values across all $n$, while uniform acceleration in Minkowski and motion in the Schwarzschild background lead to visibly larger deviations from Markovianity. Within each spacetime family, the Rindler trajectory lies above the inertial one, and the infalling worldline develops stronger deviations than the static Schwarzschild detector. Comparisons between the Minkowski and Schwarzschild families should be read as parameter dependent: different choices of $a$, $M$, or $r_{0}$ would shift the relative positions of the curves, even though the basic trend of “more structure in the correlator $\Rightarrow$ larger $N(\Upsilon)$” remains intact.
@@
Panel 5(b) reports the stitched two-slot activation $\Delta_{\text{act}}$ as a function of $n$, constructed from the excess Choi negativity of a composite two-slot block over its constituent single-slot intermediates. By design, $\Delta_{\text{act}}$ isolates superactivation-type effects: it vanishes whenever non-Markovianity is already witnessed by individual intermediates, and becomes positive only when stitching operations across well-separated diamonds unlocks additional memory resources. For the trajectories considered here, $\Delta_{\text{act}}\geq 0$ in all cases and follows the same qualitative pattern as panel 5(a): the inertial process is closest to Markovian, while accelerated and black-hole worldlines exhibit larger activation within their respective families. The initial growth and subsequent saturation of $\Delta_{\text{act}}$ with $n$ indicate that the long-range kernel contributions are converted into operational advantage up to the point where additional slots cease to probe genuinely new correlations.
@@
The scatter plots in panel 5(c) make the link between intermediate maps and the full process explicit. Here, the SDP proxy for $N(\Upsilon)$ is plotted against the certified lower bound $\max_{j<k}\tfrac{2}{d}N_{\chi}(\Phi_{k{:}j})$ obtained from the most non-CP-divisible intermediate for each geometry and $n$. The points in each panel cluster around an approximately linear trend, with best-fit lines and $R^{2}$ values close to unity. This shows that, in the weak-coupling and few-slot regime explored here, the negativity of a single “most correlated” intermediate map already captures most of the operational distance to the Markovian set. In practice, an experiment that can tomographically reconstruct only selected intermediates, rather than the full comb, still obtains a robust predictor for $N(\Upsilon)$ that is stable under changes of trajectory within a given spacetime.
@@
Panel 5(d) replaces this Choi-level quantity by a purely kernel-level measure, the normalized tail mass $S(\Delta)$ obtained by summing $|\eta_{i-j}|$ over separations $|i-j|\geq\Delta$ (here $\Delta=1$). The roughly monotone, weakly nonlinear relationship between the SDP proxy and $S(\Delta)$ across all trajectories confirms a simple “kernel-mass $\to$ operational-distance” principle: once the universal short-distance singularity has been subtracted, processes with heavier long-range tails in the regular part of the Wightman kernel tend to exhibit larger $N(\Upsilon)$. Within each family (Minkowski or Schwarzschild), this tail mass tracks the same ordering that appears in the SDP-derived bounds and in $\Delta_{\text{act}}$, indicating that the pairwise kernels distilled in Fig. 3 and Fig. 4 already contain the essential information needed to anticipate the size of the covariant non-Markovian resource.
Although our explicit numerics are performed in an effective 1+11{+}1-dimensional setting, the same qualitative picture is expected to hold in 3+13{+}1 dimensions. In that case, the Minkowski Wightman function decays faster along the worldline, which reduces the absolute size of $\eta_{i-j}$ at large separations but does not remove the hierarchy between simple inertial motion and more structured trajectories. Acceleration in flat space and the presence of a Schwarzschild horizon still introduce KMS-like periodicity and near-horizon enhancements in the pulled-back correlator, leading to longer-lived tails in the corresponding kernels. The explicit 3+13{+}1-dimensional expressions for the pulled-back Wightman functions, together with asymptotic analyses of the induced kernels and their tail masses for the trajectories considered here, are presented in the appendices and support this interpretation.
@@ Appendix B Evaluation of $\eta_{i-j}$ for smooth switching
$\chi_{\sigma}(\theta)=\frac{1}{(2\pi\sigma^{2})^{1/4}}\exp\!\left[-\frac{(\theta-\theta_{k})^{2}}{4\sigma^{2}}\right],$ (52)
centered at the midpoint $\theta_{k}$ of segment $\gamma_{k}$, with normalization $\int d\theta\,\chi_{\sigma}^{2}(\theta)=1$.
@@
$\eta_{i-j}$ $\simeq-\frac{g^{2}a^{2}}{4\pi^{2}}\int d\theta\int d\theta^{\prime}\;\chi_{\sigma}(\theta)\,\chi_{\sigma}(\theta^{\prime})\,e^{-(\theta-\theta^{\prime})}$
$\simeq-\frac{g^{2}a^{2}}{4\pi^{2}}\,e^{-\Delta\theta_{ij}}\Bigg[\int d\theta\,\chi_{\sigma}(\theta)\,e^{-\theta}\Bigg]\Bigg[\int d\theta^{\prime}\,\chi_{\sigma}(\theta^{\prime})\,e^{\theta^{\prime}}\Bigg],$ (54)
where in the second line we have expanded around the midpoints of the two windows and used the fact that the separation between midpoints is $\Delta\theta_{ij}$. For windows symmetric about their respective midpoints and of width $\sigma\lesssim\Delta\theta$, the integrals are dominated by $\theta\approx\theta_{i}$ and $\theta^{\prime}\approx\theta_{j}$ and scale as
$\int d\theta\,\chi_{\sigma}(\theta)\,e^{-\theta}\sim e^{-\theta_{i}}\sigma,\qquad\int d\theta^{\prime}\,\chi_{\sigma}(\theta^{\prime})\,e^{\theta^{\prime}}\sim e^{\theta_{j}}\sigma,$ (55)
@@
with $C_{\eta}=O(1)$ determined by the precise window shape. Thus, for smooth segment-sized windows ($\sigma\sim\Delta\theta$), the off-diagonal kernel entries decay exponentially with slot separation and scale quadratically with the effective segment width, in agreement with the qualitative behaviour used in the main text.
@@
Sharp (rectangular) switching, $\chi(\theta)=\Delta\theta^{-1/2}$ on each segment, leads to spurious high-frequency contributions and distributional boundary terms at the switching edges, as is well known in the Unruh–DeWitt literature (see, e.g., Ref. [16]). For physical predictions we therefore advocate using a smooth family $\chi_{\sigma}$ and verifying the robustness of the results either in the combined limit $\sigma\to 0$ taken after the continuum limit, or at finite $\sigma$ comparable to the slot size where observables are insensitive to moderate variations in $\sigma$.

### v2

@@ III.1 Operational Definition and Resource-Theoretic Perspectives
where the supremum ranges over causal testers (normalized multi-time instruments obeying the usual recursion $\operatorname{Tr}_{\mathrm{out}_{k}}T^{(k)}=T^{(k-1)}$ with $T^{(0)}=\mathbb{I}/d$), and the ancilla dimension can be taken no larger than the product of input dimensions up to slot n1n{-}1. Operationally, (Eq. 21) is the maximal distinguishing advantage of an optimal entanglement-assisted, adaptive probing strategy between $\Upsilon$ and its closest Markovian approximation; faithfulness follows immediately since $N(\Upsilon)=0$ iff $\Upsilon\in\mathcal{M}$ [12, 53].
@@
The distance $N$ defined in (Eq. 20, Eq. 21) has the basic features expected of an operational resource measure for memory. (i) Faithfulness and convexity: $N(\Upsilon)=0$ iff $\Upsilon\in\mathcal{M}$, and $N$ is convex on the affine space of combs: $N(p\Upsilon{+}(1{-}p)\Gamma)\leq p\,N(\Upsilon)+(1{-}p)\,N(\Gamma)$. (ii) Stability (Lipschitz continuity): by the triangle inequality for the strategy norm, $\bigl|N(\Upsilon)-N(\Gamma)\bigr|\leq\|\Upsilon-\Gamma\|_{\rm comb}$, which we use later to relate $N$ to entropic diagnostics. (iii) Monotonicity under Markovian supermaps: if $\mathbb{S}$ is a free transformation obtained by slot-wise pre/post-processing with CPTP channels and memoryless side channels that preserve the comb causality constraints (i.e., $\mathbb{S}[\mathcal{M}]\subseteq\mathcal{M}$), then
@@ III.2 Witnessing and Bounding Non-Divisibility
where $K_{k:j}=\int_{\gamma_{k:j}}\langle H_{\mathrm{int}}(\lambda)\rangle_{E}d\lambda$ (often zero) and the double commutator encodes dissipation via the Wightman function $W$. As in standard weak-coupling treatments of open quantum systems [5, 44], the second-order Born expansion provides a controlled approximation to the reduced dynamics on time scales $t\lesssim O(g^{-2})$, with any violations of complete positivity signalling use beyond its nominal regime of validity. Outside this regime one would need a nonperturbative construction of $\Upsilon$, for instance via resummation schemes or tensor-network discretizations of the field, which lies beyond the scope of the present work. The Choi operator approximates
@@ III.3 Multi-Time Phenomena and Superactivation in composite systems
Superactivation occurs whenever $\Delta_{\mathrm{act}}>0$, indicating memory effects undetectable in marginal probes but revealed by coherent multi-time interventions (cross-time links in Figure 2(b)). From the perspective of relativistic quantum information, these multi-probe effects are complementary to entanglement-harvesting protocols in which UDW pairs act as curvature and horizon probes via their shared vacuum correlations [21, 34], but here the emphasis is on temporal structure and a fully multi-time operational distance.
@@
where cross-segment kernels $\{\eta_{m-l}\}_{l\in S_{1},m\in S_{2}}$ add coherently. Curvature-induced long-range tails in $W$ (e.g., near horizons) can make intra-segment sums positive (CP maps) while cross terms drive overall negativity, yielding $\Delta_{\mathrm{act}}>0$ and a finite LB (Eq. (39)) [42, 26, 51, 33, 41].
@@ IV Memory kernel structures in curved spacetimes: numerical benchmarks and physical insights
To isolate curvature-induced memory, we employ Hadamard subtraction [27, 16, 25]: the universal singular part of $W$ (arising from the Minkowski-like local structure near coincidence) is removed, leaving the regular, state-dependent remainder that encodes global correlations. This subtraction ensures that $\eta$ reflects genuine non-local effects rather than ultraviolet divergences. Numerically, we discretize the worldline into $n=14$ slots over a total affine parameter range $L=12.0$, with Gaussian switches of width $w=0.4\,\Delta$ (where $\Delta=L/n$) and use Simpson quadrature with 500 points per integral for accuracy. Geodesic integration for infalling trajectories is handled via solve_ivp with adaptive Runge-Kutta (RK45) stepping, achieving relative tolerances of $10^{-8}$.
@@
Figure 4 displays line cuts of the memory kernel against slot separation together with parameter scans of the kernel’s most negative entry, which we use as a proxy for long–range anticorrelations. Panel 4(a) plots the median magnitude $(\langle|\eta_{i-j}|\rangle)$ versus (|ij|)(|i-j|) on semi–log axes for four worldlines: Minkowski inertial (blue), Minkowski Rindler (orange), Schwarzschild static (green), and Schwarzschild infall (red). Across all scenarios except Schwarzschild–static, we observe a dip at |ij|=1|i-j|=1, which arises because adjacent slots correspond to partially overlapping causal diamonds: the convolution of the Wightman kernel with two overlapping window functions produces strong destructive cancellation. As the separation grows, the diamonds cease to overla;, this cancellation rapidly weakens, and the median $\langle|\eta_{i-j}|\rangle$ rebounds toward the asymptotic floor set by the intrinsic tail of the correlator (algebraic for inertial/infall, thermal for Rindler/static). This dip is therefore best viewed as a feature of the discretization rather than a universal short-distance signature of memory. The Schwarzschild–static case lacks this feature because proper-time stationarity with a single Tolman thermal scale places the kernel directly on its exponential slope, so the overlap-induced minimum is washed out.
@@
Figure 4(b)–(d) quantify how geometry and kinematics tune the most negative off–diagonal entry. For infall, $(\min\eta)$ becomes more negative as the initial radius $(r_{0})$ increases [Figure 4(b)], indicating that longer exposure to gentle curvature gradients amplifies the negative lobe at intermediate lags. In the Rindler case, $(\min\eta)$ plateaus for small accelerations and then decreases approximately linearly beyond a knee at $(a\Delta\tau=\mathcal{O}(1))$ [Figure 4(c)]: below the knee the Unruh correlation time $(\xi_{U}\propto 1/a)$ exceeds the slot width and the kernel is effectively inertial; once resolved, thermal anticorrelations deepen with (a). For static Schwarzschild, $(\min\eta)$ shows a shallow shoulder at small radii before a gentle outward decrease [Figure 4(d)]. This shape reflects the competition between Tolman heating, which shortens $(\xi(r))$ and strengthens anticorrelations near the horizon, and redshift of the proper-time step $(\Delta\tau=\sqrt{1-2M/r},\Delta t)$ at fixed coordinate sampling, which reduces adjacent–slot overlap and partially offsets thermal screening. Far from the horizon the local temperature drops, the correlation length grows, and $(\min\eta)$ relaxes toward the cooler, nearly flat–space limit.
@@
Figure 5 gathers these kernel-level observations into fully operational statements about the process tensor. Figure 5(a) shows, for $n\leq 5$, the SDP-derived lower-bound proxy for $N(\Upsilon)$ obtained by fitting the tester data to the nearest CP-divisible product of single-step channels (see D. For $n\leq 5$ we have performed the full SDP on a single workstation; however, the exponential growth of the underlying Hilbert-space dimension makes such optimizations rapidly intractable for larger $n$. The bound is quantified by the summed absolute prediction error $\sum_{\text{testers}}|\Delta|$, which lower-bounds the strategy-norm distance between the true process tensor and its best Markovian approximation within this restricted class. For our benchmark parameters, the Minkowski inertial worldline yields the smallest values across all $n$, while uniform acceleration in Minkowski and motion in the Schwarzschild background lead to visibly larger deviations from Markovianity. Within each spacetime family, the Rindler trajectory lies above the inertial one, and the infalling worldline develops stronger deviations than the static Schwarzschild detector. Comparisons between the Minkowski and Schwarzschild families should be read as parameter dependent: different choices of $a$, $M$, or $r_{0}$ would shift the relative positions of the curves, even though the basic trend of “more structure in the correlator $\Rightarrow$ larger $N(\Upsilon)$” remains intact.
@@
Figure 5(b) reports the stitched two-slot activation $\Delta_{\text{act}}$ as a function of $n$, constructed from the excess Choi negativity of a composite two-slot block over its constituent single-slot intermediates. By design, $\Delta_{\text{act}}$ isolates superactivation-type effects: it vanishes whenever non-Markovianity is already witnessed by individual intermediates, and becomes positive only when stitching operations across well-separated diamonds unlocks additional memory resources. For the trajectories considered here, $\Delta_{\text{act}}\geq 0$ in all cases and follows the same qualitative pattern as figure 5(a): the inertial process is closest to Markovian, while accelerated and black-hole worldlines exhibit larger activation within their respective families. The initial growth and subsequent saturation of $\Delta_{\text{act}}$ with $n$ indicate that the long-range kernel contributions are converted into operational advantage up to the point where additional slots cease to probe genuinely new correlations.
@@
The scatter plots in Figure 5(c) make the link between intermediate maps and the full process explicit. Here, the SDP proxy for $N(\Upsilon)$ is plotted against the certified lower bound $\max_{j<k}\tfrac{2}{d}N_{\chi}(\Phi_{k{:}j})$ obtained from the most non-CP-divisible intermediate for each geometry and $n$. The points in each panel cluster around an approximately linear trend, with best-fit lines and $R^{2}$ values close to unity. This shows that, in the weak-coupling and few-slot regime explored here, the negativity of a single “most correlated” intermediate map already captures most of the operational distance to the Markovian set. In practice, an experiment that can tomographically reconstruct only selected intermediates, rather than the full comb, still obtains a robust predictor for $N(\Upsilon)$ that is stable under changes of trajectory within a given spacetime.
@@
Figure 5(d) replaces this Choi-level quantity by a purely kernel-level measure, the normalized tail mass $S(\Delta)$ obtained by summing $|\eta_{i-j}|$ over separations $|i-j|\geq\Delta$ (here $\Delta=1$). The roughly monotone, weakly nonlinear relationship between the SDP proxy and $S(\Delta)$ across all trajectories confirms a simple “kernel-mass $\to$ operational-distance” principle: once the universal short-distance singularity has been subtracted, processes with heavier long-range tails in the regular part of the Wightman kernel tend to exhibit larger $N(\Upsilon)$. Within each family (Minkowski or Schwarzschild), this tail mass tracks the same ordering that appears in the SDP-derived bounds and in $\Delta_{\text{act}}$, indicating that the pairwise kernels distilled in Figure 3 and Figure 4 already contain the essential information needed to anticipate the size of the covariant non-Markovian resource.
Although our explicit numerics are carried out in an effective $(1+1)$-dimensional setting, the qualitative hierarchy between trajectories is expected to persist in $(3+1)$ dimensions with a smaller overall scale of the memory effects. In $(3+1)$-dimensional Minkowski spacetime the vacuum Wightman function along a timelike worldline decays as $1/s^{2}$, whereas in $(1+1)$ dimensions it exhibits a logarithmic dependence (Eq. 43). After smearing with smooth windows of fixed macroscopic size, this faster decay reduces the absolute magnitude of the off-diagonal kernel entries $\eta_{i-j}$ and hence the tail mass $S(\Delta)$, but does not alter the ordering between inertial, accelerated, and near-horizon trajectories. Acceleration and the presence of a Schwarzschild horizon still imprint KMS-like periodicity and near-horizon enhancements in the pulled-back correlator, which translate into parametrically longer-lived kernel tails than in flat inertial motion. Since both the Choi-negativity lower bound and the SDP-based proxy for $N(\Upsilon)$ depend linearly on these tails, the operational advantage in $(3+1)$ dimensions is suppressed only by an order-one factor set by microscopic versus slot scales, rather than exponentially in the number of dimensions. Analytical estimates for the $(3+1)$-dimensional kernels for the same trajectories, presented in the appendices, support this expectation.
@@ V Conclusion
DATA AVAILABILITY
The numerical data and analysis scripts that support the findings of this work are openly available in the Zenodo repository [52].
@@ Appendix B Evaluation of $\eta_{i-j}$ for smooth switching
$\chi_{w}(\theta)=\frac{1}{(2\pi w^{2})^{1/4}}\exp\!\left[-\frac{(\theta-\theta_{k})^{2}}{4w^{2}}\right],$ (52)
centered at the midpoint $\theta_{k}$ of segment $\gamma_{k}$, with normalization $\int d\theta\,\chi_{w}^{2}(\theta)=1$.
@@
$\eta_{i-j}$ $\simeq-\frac{g^{2}a^{2}}{4\pi^{2}}\int d\theta\int d\theta^{\prime}\;\chi_{w}(\theta)\,\chi_{w}(\theta^{\prime})\,e^{-(\theta-\theta^{\prime})}$
$\simeq-\frac{g^{2}a^{2}}{4\pi^{2}}\,e^{-\Delta\theta_{ij}}\Bigg[\int d\theta\,\chi_{w}(\theta)\,e^{-\theta}\Bigg]\Bigg[\int d\theta^{\prime}\,\chi_{w}(\theta^{\prime})\,e^{\theta^{\prime}}\Bigg],$ (54)
where in the second line we have expanded around the midpoints of the two windows and used the fact that the separation between midpoints is $\Delta\theta_{ij}$. For windows symmetric about their respective midpoints and of width $w\lesssim\Delta\theta$, the integrals are dominated by $\theta\approx\theta_{i}$ and $\theta^{\prime}\approx\theta_{j}$ and scale as
$\int d\theta\,\chi_{w}(\theta)\,e^{-\theta}\sim e^{-\theta_{i}}w,\qquad\int d\theta^{\prime}\,\chi_{w}(\theta^{\prime})\,e^{\theta^{\prime}}\sim e^{\theta_{j}}w,$ (55)
@@
with $C_{\eta}=O(1)$ determined by the precise window shape. Thus, for smooth segment-sized windows ($w\sim\Delta\theta$), the off-diagonal kernel entries decay exponentially with slot separation and scale quadratically with the effective segment width, in agreement with the qualitative behaviour used in the main text.
@@
Sharp (rectangular) switching, $\chi(\theta)=\Delta\theta^{-1/2}$ on each segment, leads to spurious high-frequency contributions and distributional boundary terms at the switching edges, as is well known in the Unruh-DeWitt literature (see, e.g., Ref. [33]). For physical predictions we therefore advocate using a smooth family $\chi_{\sigma}$ and verifying the robustness of the results either in the combined limit $w\to 0$ taken after the continuum limit, or at finite $w$ comparable to the slot size where observables are insensitive to moderate variations in $w$.